Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 563
Entry time: 06 Nov 2024, 09:38
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 06.11.2024 |
|
|
General Announcement:
- Invitation for scheduling meeting was distributed to some groups. The Winter maintenance period is relatively short.
- UOM next Friday takes place with presentations from FD studies and jitter reports from the instruments.
CTRL:
- Nothing to report.
DA:
- Nothing to report.
EEE:
- Nothing to report.
VAC:
- Nothing to report; all fine.
XPD:
- Nothing to report; no news on PES.
DET:
- Nothing to report.
- Successful tests with high Z yesterday.
ITDM:
- Nothing to report.
XRO:
- Nothing major to report; prepared for the 30keV studies.
LAS:
- Nothing to report.
UNSYS:
- Yesterday problem with PS motor in cell 25/SA1: solved by ZZ;
- Apple-X and PS cell 26/SA3: tomorrow 7AM a ZZ to try to fix problems.
SPB:
- All fine.
FXE:
- Not present.
MID:
- Not present.
HED:
- Not present.
SCS:
- Nothing to report.
SQS:
- Nothing to report.
SXP:
- Not present. |